# Privacy-preserving fall detection at the edge using Sony IMX636 event-based vision sensor and Intel Loihi 2 neuromorphic processor

### v1

@@ V-A Improving fall detection score
While further algorithmic improvements are likely necessary, a promising approach involves learning the variance of predictions in a self-supervised way. This enables the modeling of input-dependent aleatoric uncertainty [Kendall_Yarin17], which can encourage more robust and reliable predictions [paredes-valles_etal25].
@@ V-B Reducing power consumption
Second, at the algorithmic level, we have seen that activation sparsity is the highest for stateful spiking neurons, where the state is either used specifically to sparsify activations as in SigmaDelta neurons, or to further add temporal feature extraction as in LIF neurons. Stateful models are the most promising for achieving the highest sparsity while maintaining a high accuracy because they can exploit the context encoded in the states [Zhou_etal25], at the cost of a higher memory footprint. While stateful processing cannot be applied to the patched MCU13B since Loihi 2 memories are overwritten for each patch, it can be applied to the CNN+S4D model where a variant of the SigmaDelta neuron can be used to sparsify the SSM neurons activations.
@@ V-C On-chip anonymization of EVS data
Fully embedded processing is necessary for privacy-preserving fall detection, but it may not be sufficient at the application level as some EVS data may have to be be sent to a cloud service in the case of a fall detection, for example, to confirm the need for medical assistance. However, raw events should not be transmitted, because even without full RGB content they can be used to reconstruct grayscale frames, which undermines privacy protection. [Rebecq_etal21, Paredes-Valles_etal21, Zhu_etal22]. Therefore, on-chip anonymization of the event data that would be transferred (e.g., a sliding-window buffer of a few seconds) is required in an always-on fashion, in parallel to the fall detection inference.
For example, authors in [Ahmad_etal23] try to anonymize event streams to protect the identity of human subjects against image reconstruction, by scrambling events and enforcing the degradation of such images. However, deploying a similar algorithm in any embedded device remains challenging in terms of required computing resources and power. Further research is required to get a fully privacy-preserving ecosystem which will enable the adoption of security cameras in sensitive environments, using neuromorphic sensing and processing in particular.
@@ VI Conclusion
We showed that the proposed graded spikes in LIF models increase the score by 6.2% and decrease the number of SynOps/s by 4.8 compared to binary spikes, reaching an overall SynOps/s sparsity of 55.5 compared to an iso-topology neural network with dense processing. Compared to the binary LIF, the graded LIF comes at the cost of an extra multiplication for each SynOp which is nevertheless natively supported by Loihi 2 thanks to its programmable neuro-core [Shrestha_etal24]. In addition, it was shown that the main advantage of SNN over ANN accelerators comes from exploiting the sparsity of spikes and not from the replacement of MAC by AC operations [Dampfhoffer_etal23]. Moving forward, graded spikes can expand the design space of neuromorphic computing algorithms beyond binary SNNs toward a broader category of sparse neural networks, including but not limited to graded LIF and SigmaDelta neurons. Further research on hardware and algorithms is required to explore this new paradigm of graded spikes, which can also refine the current abstraction level from biological neurons, moving beyond the binary spike representation to focus on sparsity.
@@
Furthermore, we showed for the first time that we can deploy a SOTA MCU13B spatial feature extractor on neuromorphic hardware, using patched inference to fit the model in a single Loihi 2 chip. Coupled with the S4D temporal feature extractor, the model achieves the highest score of 83.6 %, satisfies our real-time constraint with 16 predictions/s and only consumes 90 mW on Loihi 2. Nevertheless, about 90% of the total power is static power, which is the main limitation of the hardware system. Such static power can be reduced by removing unused components and manufacturing the next version of the Loihi chip with a more advanced technology process. Our system serves as a proof-of-concept demonstrating the potential of embedding dedicated processing at the edge of EVS via a direct interface, in which events are represented as graded spikes and processed directly to fully leverage their spatio-temporal sparsity. This approach enables low-latency and low-power perception at the system level.
@@
Beyond the privacy-preserving aspect of edge computing, the always-on perception serves as a filter of raw sensory data by only sending out some data when a pattern of interest is detected (i.e., something relevant happens). This drastically reduces the amount of data that need to be transferred, processed and stored in the cloud and thus reduces the overall energy consumption of the application ecosystem. Neuromorphic technologies can enable this transition from cloud-centric to edge-distributed real-time processing toward a more balanced, efficient and scalable ecosystem, while leveraging the growth of IoT and edge sensing markets [YOLE_24, Muir_Sheik25]. By starting from the application requirements and applying a system-level optimization through coherent sensor–processor–algorithm co-development, this work demonstrates a concrete step toward bringing neuromorphic technologies into the real world.

### v2

@@ V-B Reducing power consumption
Second, at the algorithmic level, we have seen that activation sparsity is the highest for stateful spiking neurons, where the state is either used specifically to sparsify activations as in SigmaDelta neurons, or to further add temporal feature extraction as in LIF neurons. Stateful models are the most promising for achieving the highest sparsity while maintaining a high accuracy because they can exploit the context encoded in the states [63], at the cost of a higher memory footprint. While stateful processing cannot be applied to the patched MCU13B since Loihi 2 memories are overwritten for each patch, it can be applied to the CNN+S4D model where a variant of the SigmaDelta neuron can be used to sparsify the SSM neurons activations.
@@ V-C On-chip anonymization of EVS data
Fully embedded processing is necessary for privacy-preserving fall detection, but it may not be sufficient at the application level as some EVS data may have to be be sent to a cloud service in the case of a fall detection, for example, to confirm the need for medical assistance. However, raw events should not be transmitted, because even without full RGB content they can be used to reconstruct grayscale frames, which undermines privacy protection. [49, 44, 64]. Therefore, on-chip anonymization of the event data that would be transferred (e.g., a sliding-window buffer of a few seconds) is required in an always-on fashion, in parallel to the fall detection inference.
For example, authors in [1] try to anonymize event streams to protect the identity of human subjects against image reconstruction, by scrambling events and enforcing the degradation of such images. However, deploying a similar algorithm in any embedded device remains challenging in terms of required computing resources and power. Further research is required to get a fully privacy-preserving ecosystem which will enable the adoption of security cameras in sensitive environments, using neuromorphic sensing and processing in particular.
@@ VI Conclusion
We showed that the proposed graded spikes in LIF models increase the score by 6.2% and decrease the number of SynOps/s by 4.8 compared to binary spikes, reaching an overall SynOps/s sparsity of 55.5 compared to an iso-topology neural network with dense processing. Compared to the binary LIF, the graded LIF comes at the cost of an extra multiplication for each SynOp which is nevertheless natively supported by Loihi 2 thanks to its programmable neuro-core [56]. In addition, it was shown that the main advantage of SNN over ANN accelerators comes from exploiting the sparsity of spikes and not from the replacement of MAC by AC operations [17]. Moving forward, graded spikes can expand the design space of neuromorphic computing algorithms beyond binary SNNs toward a broader category of sparse neural networks, including but not limited to graded LIF and SigmaDelta neurons. Further research on hardware and algorithms is required to explore this new paradigm of graded spikes, which can also refine the current abstraction level from biological neurons, moving beyond the binary spike representation to focus on sparsity.
@@
Furthermore, we showed for the first time that we can deploy a SOTA MCU13B spatial feature extractor on neuromorphic hardware, using patched inference to fit the model in a single Loihi 2 chip. Coupled with the S4D temporal feature extractor, the model achieves the highest score of 83.6 %, satisfies our real-time constraint with 16 predictions/s and only consumes 90 mW on Loihi 2. Nevertheless, about 90% of the total power is static power, which is the main limitation of the hardware system. Such static power can be reduced by removing unused components and manufacturing the next version of the Loihi chip with a more advanced technology process. Our system serves as a proof-of-concept demonstrating the potential of embedding dedicated processing at the edge of EVS via a direct interface, in which events are encoded as graded spikes and processed directly to fully leverage their spatio-temporal sparsity. This approach enables low-latency and low-power perception at the system level.
@@
Beyond the privacy-preserving aspect of edge computing, the always-on perception serves as a filter of raw sensory data by only sending out some data when a pattern of interest is detected (i.e., something relevant happens). This drastically reduces the amount of data that need to be transferred, processed and stored in the cloud and thus reduces the overall energy consumption of the application ecosystem. Neuromorphic technologies can enable this transition from cloud-centric to edge-distributed real-time processing toward a more balanced, efficient and scalable ecosystem, while leveraging the growth of IoT and edge sensing markets [28, 40]. By starting from the application requirements and applying a system-level optimization through coherent sensor–processor–algorithm co-development, this work demonstrates a concrete step toward bringing neuromorphic technologies into the real world.